# CLAPSep: Leveraging Contrastive Pre-trained Model for Multi-Modal Query-Conditioned Target Sound Extraction

### v3

@@ I Introduction
Note that, most existing query-conditioned methods only consider positive queries to indicate “what sound to extract.” In our study, we discover that providing explicit information to the model about “what sound to suppress” further aids in target sound extraction. Therefore, we design a strategy to facilitate the model to support both positive queries and negative queries to perform target sound extraction and target sound removal (TSR) in a unified model. This design can not only enhance the model flexibility, but also lead to enhanced model performance when both positive and negative queries are provided.
@@
    We introduce CLAPSep, a CLAP-based target sound extraction model to perform query-conditioned target sound extraction in real-world sound mixtures. By reusing the pre-trained CLAP encoders to form a tightly coupled system, based on which parameter-efficient modules are designed to fit specific downstream tasks of TSE, we train a powerful query-conditioned TSE model both data-efficiently and with reduced computational demands;
@@
    We provide extensive experiments to compare our method with previous benchmarks. Experimental results demonstrate that our method achieves state-of-the-art (SOTA) performance in target sound extraction. The results of zero-shot experiments indicate that our proposed CLAPSep has a good model generalizability. We also carry out ablation studies to demonstrate the effectiveness of each individual component within our design.
@@ II-C Query-Conditioned Target Sound Extraction
Query-conditioned target sound extraction offers an alternative approach to dealing with sound mixtures by focusing solely on extracting the desired sound while disregarding all other sources present in real-world audio mixtures. In the realm of query-conditioned target sound extraction, the separation model’s operation is conditioned on a query that specifies the particular sound to be extracted from the mixture. Existing methods for query-conditioned target sound extraction can be classified into four categories based on different types of queries: label-querying methods [13, 14, 15], visual-querying methods [18, 19], audio-querying methods [9, 15, 16, 17], and language-querying methods [20, 21, 16].
@@
Flexible adaptation to different types of user queries is also a very important point and directly determines the user-friendliness of the query-conditioned TSE systems. To support multimodal queries, prior work [33] proposes a cross-attention module to aggregate textual, visual, and label-formed queries and [15] incorporate a mixed-encoder mechanism to support one-hot label and audio queries for targeted sound extraction. In [34], researchers propose to tune a large language model (LLM) with LoRA to complement the traditional voice print with text for target speech extraction. While these efforts enhance the flexibility of TSE through multimodal queries, the query networks also have to train from scratch together with the separation network to comprehend audio events, which would lead to sub-optimal generalizability for unseen sound events during model training. Moreover, these methods are limited in flexibility as they only support positive queries and cannot perform the target sound removal task. In [35], authors propose to LoRA-tune an LLM as a query network to perform language-queried target sound extraction and target sound removal in a unified model. This is achieved by explicitly adding the keywords “extract” or “remove” to the language query. Although this approach has a user-friendly form of interaction, the introduction of an LLM imposes a significant computational overhead and fails to support multimodal queries. Furthermore, due to the language model’s insufficient attention to negative keywords, the model in [35] shows a significant performance gap between the TSE and TSR tasks. Appropriate incorporation of multimodal and multi-polarity queries for flexible and generalizable TSE and TSR in a unified model is still under-explored.
@@ II-D Leveraging Prior Knowledge in Pre-Trained Models
The success of multi-modal contrastive pre-training [26, 23] makes it possible to decouple the training of the query model and separation model in query-conditioned target sound extraction systems, thus facilitating the use of prior knowledge embedded in the pre-trained models for better model generalizability in the downstream task of TSE. The CLAP model is composed of an audio encoder and a text encoder. In the pre-training phase of CLAP, the model was trained with the contrastive learning paradigm between the audio-text pairs, bringing audio and text descriptions into a joint multi-modal embedding space. Thus, the encoded language query embeddings can be used to guide the separation model to indicate what sound to extract. In [18], a query-conditioned TSE model called CLIPSep was proposed, which employed the CLIP text and image encoder to extract language and vision embeddings, serving as queries for TSE. Importantly, the query models remained frozen during the training of the separation model, which helps maintain the semantic space of the pre-trained model and aids in model convergence and generalizability. In AudioSep [21], researchers also leveraged the pre-trained text encoder of CLIP and CLAP as the language querying model. However, a notable limitation in these approaches is that while pre-trained models are appropriately used as query networks, the separation networks are randomly initialized which can not establish a prior semantic association with the query networks to comprehend sound events. As a result, these models still necessitate a substantial amount of data and computational resources for better model performance and generalizability.
@@ III-B5 Loss Function
$\mathrm{SISDR}(\mathbf{\hat{x}},\mathbf{x})$ $=10\mathrm{log}_{10}\left(\frac{||\frac{||\mathbf{\hat{x}}^{T}\mathbf{x}||^{2}}{||\mathbf{x}||^{2}}\mathbf{x}||^{2}}{||\frac{||\mathbf{\hat{x}}^{T}\mathbf{x}||^{2}}{||\mathbf{x}||^{2}}\mathbf{x}-\mathbf{\hat{x}}||^{2}}\right),$ (17)
@@ IV-A2 Test Data
We use multiple datasets to perform a comprehensive evaluation of our proposed method. This includes AudioCaps [42], AudioSet [43], ESC-50 [45] and FSDKaggle2018 [46] for universal sound separation and MUSIC21 [19] for musical instrument separation. In each dataset, evaluation sound mixtures are generated by considering each audio clip as a target source. Subsequently, either 1 or 5 additional audio clips are randomly chosen as interference noise. The mixing process involves combining one target source with one interference noise at a signal-to-noise ratio (SNR) of 0dB, resulting in the creation of 1 or 5 evaluation mixtures for each target source. For label-annotated datasets, we transform the labels into language descriptions by adding the prefix “The sound of ”. For ESC-50, FSDKaggle2018, and MUSIC21, we also randomly choose 10 audio clips for each sound class as query audio samples to evaluate the multi-modal clues queried TSE capability of the proposed approach. These selected query samples are not used to generate evaluation mixtures, in order to avoid information leakage. In the following section, we will introduce all the evaluation datasets used in our experiments.
@@
TABLE II: Language-queried target sound extraction performance evaluation and SOTA comparison.
@@
ESC-50 is a labeled collection of environmental audio clips extracted from public field recordings gathered by the Freesound project [47]. The ESC-50 dataset comprises 2,000 5-second-long audio clips belonging to 50 environmental sound classes such as car horn, and pouring water, among others. Each audio clip is assigned to one class label and all the audio clips are sampled at 44.1kHz. For consistency comparison, we downsample the audio clips to 32kHz. Since there is no official test split, we use the whole dataset to create 6,500 evaluation sound mixtures after filtering 500 audio clips used as query audio samples.
FSDKaggle2018 consists of 11,073 audio files annotated with 41 labels, covering both natural sounds and instrumental music. All the sound samples are gathered from the Freesound project [47] and sampled at 44.1kHz. We resample the audio clips to 32kHz for consistency. To create evaluation sound mixtures, we use the official test split which contains 1,600 audio clips to create 8,000 evaluation sound mixtures.
@@ V-A Language-Queried TSE Performance Evaluation
TABLE III: multi-modal clues queried target sound extraction performance evaluation and SOTA comparison.
@@
The proposed CLAPSep model consistently outperforms other models by a significant margin, except for the MUSIC21 dataset. This anomaly can be traced back to the composition of the AudioCap dataset, on which the CLAPSep trained, contains only 169 music clips among its total of 49,274 clips. Impressively, despite being trained on a considerably smaller dataset than AudioSep, CLAPSep exhibits superior generalizability. This outcome underscores our model’s proficiency in utilizing the extensive sound category knowledge encoded in the pre-trained CLAP model, achieved through integrating the CLAP audio encoder into the separation network. It is worth noting that CLAPSep supports both positive and negative queries. Its performance with negative prompts is comparable to or even better than that with positive prompts. This is because we treat positive and negative queries equally in constructing the condition embedding so that the user will not suffer from the performance gap between TSE and TSR, which is significant in [35] in contrast. Furthermore, the two types of prompts are found to be complementary in the process of extracting target sounds. When CLAPSep is used with both positive and negative prompts, it demonstrates significantly improved performance, particularly on the MUSIC21 dataset. This indicates that models like CLAPSep, which accommodate fine-grained queries, can leverage more complex input queries to effectively extract target sounds, even from sound categories that are rarely represented in the training data.
Fine-grained queries improve not only the sound extraction performance in terms of SDRi and SISDRi but also the consistency and stability among different conditions. We visualize the SNRi and SISINRi distribution of CLAPSep and AudioSep across the 5 datasets, as shown in Fig. 4. CLAPSep has higher median scores and is more centrally distributed with lower variance, indicating its better performance and stability.
@@
V-B Multi-Modal Clues Queried TSE Performance Evaluation
TABLE IV: Model performance varying with different numbers of sound sources.
@@ V-B Multi-Modal Clues Queried TSE Performance Evaluation
For comparison, we employ the representative audio-queried TSE model named Universal Source Separation (USS) [9]. Notably, USS is trained on an extensively large-scale dataset, in comparison to ours. The results are presented in Table III, where “Shots” denotes the number of query audio clips chosen for each sound class. Note that the input query audio clips are converted into fixed-dimensional audio embeddings. When a query has multiple audio samples for a target sound event, the audio embeddings are averaged as one single conditional embedding.
Table III presents the evaluation results with multi-modal queries. Provided with audio-only queries, the CLAPSep-text model (trained with text-only queries) performs significantly worse than all the others. This underperformance can be attributed to the modality gap phenomenon [37]. Despite the robust multi-modal alignment capabilities of the CLAP encoders, the modality gap persistently exists and substantially hampers the separation decoder trained on a single modality to handle multi-modal queries. In contrast, the model trained with multi-modal queries (denoted as CLAPSep-hybrid) demonstrates remarkable and consistent improvements, affirming the effectiveness of the multi-modal training strategy (as indicated in equation 3) in mitigating the modality gap between CLAP audio and text encoders. Compared with USS, the proposed model achieves state-of-the-art performance on ESC-50 and FSDKaggle2018 when queried by positive queries only. In the circumstance where negative queries are available, CLAPSep outperforms USS by more than 1 dB on MUSIC21. The query audio embedding can also be augmented with text embedding (referred to as “Audio+Text”). In this case, the proposed CLAPSep model achieves the best performance on most benchmarks. This underscores the effectiveness of our proposed method in leveraging clues from multi-modal queries for target sound extraction.
Fig. 5 analyses how the number and duration of query audio samples affect the performance. The number of query audio samples, referred to as “Shots,” plays a significant role in audio-queried TSE performance. When audio samples are selected randomly, a higher number of samples offers richer semantic information, thereby enabling a more accurate representation of the sound class. Conversely, the duration of query audio samples exhibits a “sweet spot” effect. Specifically, when the duration of query audio samples reaches 3 seconds, the TSE system demonstrates its optimal performance. This phenomenon suggests that a 3-second clip of query audio can adequately capture the common characteristics of most sound events. Longer duration carves out more detailed but irrelevant characteristics associated with the audio clip that can potentially diminish performance.
@@ V-C Zero-Shot Generalizability Evaluation
In this section, we assess the zero-shot capacities of CLAPSep on source sound events that are not seen during training. The AudioSet_balanced_train set is employed due to its well-defined classification of event types. Following the setting in [49], we select 10 sound classes of AudioSet as unseen and exclude all audio clips corresponding to these classes from the training set. To generate evaluation sound mixtures, we randomly select 50 audio clips per sound class in AudioSet_test, resulting in 500 source audio clips in total for evaluation. Each audio clip, regarded as the target source, is mixed with 1 randomly selected audio clip and repeated 5 times to create 5 audio mixtures. The mixture is made at a signal-to-noise (SNR) ratio of 0 dB. To this end, we produced 250 sound mixtures per sound class, resulting in a total of 2,500 sound mixtures for evaluation.
The results are shown in Fig. 6 and denoted as “Unseen”. For comparison, we also train CLAPSep without filtering the audio event classes, denoted “Seen”. To mitigate potential label leakage, only positive queries are used for evaluation. The performance gap between “Unseen” and “Seen” is generally minor, indicating that CLAPSep can effectively generalize to sound event classes that are unseen during its training phase. It means CLAPSep can inherit the prior knowledge encoded in the CLAP audio encoder to identify and extract diverse sound events.
V-D Practical Scenarios Analysis and Evaluation
TABLE V: Model performance with input SDR varying from -10 to 10 dB.
@@ V-D Practical Scenarios Analysis and Evaluation
TABLE VI: Ablation study of language-queried TSE performance.

### v5

@@ I Introduction
Note that, most existing query-conditioned methods only consider positive queries to indicate “what sound to extract.” In our study, we discover that providing explicit information to the model about “what sound to suppress” further aids in target sound extraction. Therefore, we design a strategy to facilitate the model to support both positive queries and negative queries to perform target sound extraction and target sound removal (TSR) in a unified model. This design can not only enhance the model flexibility but also lead to enhanced model performance when both positive and negative queries are provided.
@@
    We introduce CLAPSep, a CLAP-based target sound extraction model to perform query-conditioned target sound extraction in real-world sound mixtures. By reusing pre-trained CLAP encoders both in query and separation modules to form a tightly coupled system, based on which light-weight modules are designed to fit specific downstream tasks of TSE, we train a powerful query-conditioned TSE model both data-efficiently and with reduced computational demands;
@@
    We provide extensive experiments to compare our method with previous state-of-the-art (SOTA) methods. Experimental results demonstrate that our method achieves new SOTA performance in target sound extraction. Zero-shot evaluations demonstrate the superior generalizability of the proposed method. We also carry out ablation studies to demonstrate the effectiveness of each individual component within our design.
@@ II-C Query-Conditioned Target Sound Extraction
Query-conditioned target sound extraction offers an alternative approach to dealing with sound mixtures by focusing solely on extracting the desired sound while disregarding all other sources present in real-world audio mixtures. In the realm of query-conditioned target sound extraction, the separation model’s operation is conditioned on a query that specifies the particular sound to be extracted from the mixture. Existing methods for query-conditioned target sound extraction can be classified into four categories based on different types of queries: label-queried methods [13, 14, 15], visual-queried methods [18, 19], audio-queried methods [9, 15, 16, 17], and language-queried methods [20, 21, 16].
@@
Flexible adaptation to different types of user queries is also a very important point and directly determines the user-friendliness of the query-conditioned TSE systems. To support multi-modal queries, prior work [33] proposes a cross-attention module to aggregate textual, visual, and label-formed queries and [15] incorporate a mixed-encoder mechanism to support one-hot label and audio queries for targeted sound extraction. Although these efforts improve the flexibility of TSE models through multi-modal queries, these query networks still need to be trained alongside the separation networks to understand sound events, which can lead to sub-optimal generalizability for sound events that were not encountered during the training phase. Moreover, these methods are limited in flexibility as they only support positive queries and cannot perform the target sound removal task. Recent works [34, 35] attempt to tune LLM for flexible querying of target speech/sound extraction models. In [35], authors propose to LoRA-tune an LLM as a query network to perform language-queried target sound extraction and target sound removal in a unified model. This is achieved by explicitly adding the keywords “extract” or “remove” to the language query. Although this approach has a user-friendly form of interaction, the introduction of an LLM imposes a significant computational overhead and fails to support multi-modal queries. Furthermore, due to the language model’s insufficient attention to negative keywords, the model in [35] shows a significant performance gap between the TSE and TSR tasks. Appropriate incorporation of multi-modal and multi-polarity queries for flexible and generalizable TSE and TSR in a unified model is still under-explored.
@@ II-D Leveraging Prior Knowledge in Pre-Trained Models
The success of contrastive language-audio pre-training [23] has unified audio and text into a joint embedding space. Leveraging the prior knowledge embedded in this multi-modal embedding space has been shown to enhance a range of audio-language downstream tasks, such as text-to-audio generation [36, 37, 38], audio captioning [39, 40], and more. There are also some successful practices in the task of target sound extraction. In [18], a query-conditioned TSE model called CLIPSep was proposed, which employed the CLIP text and image encoder to extract language and vision embeddings, serving as queries for TSE. Importantly, the query models remained frozen during the training of the separation model, which helps maintain the semantic space of the pre-trained model and aids in model generalizability. In AudioSep [21], researchers also leveraged the pre-trained text encoder of CLIP and CLAP as the language querying model. However, a notable limitation in these approaches is that while pre-trained models are appropriately used as query networks, the separation networks are randomly initialized which can not establish a prior semantic association with the query networks to comprehend sound events. As a result, these models still necessitate a substantial amount of data and computational resources for better model performance and generalizability.
@@ III-B5 Loss Function
$\mathrm{SISDR}(\mathbf{\hat{x}},\mathbf{x})$ $=10\mathrm{log}_{10}\left(\frac{||\frac{\mathbf{\hat{x}}^{\top}\mathbf{x}}{||\mathbf{x}||^{2}}\mathbf{x}||^{2}}{||\frac{\mathbf{\hat{x}}^{\top}\mathbf{x}}{||\mathbf{x}||^{2}}\mathbf{x}-\mathbf{\hat{x}}||^{2}}\right),$ (17)
@@ IV-A2 Test Data
We use multiple datasets to perform a comprehensive evaluation of our proposed method. This includes AudioCaps [47], AudioSet [48], ESC-50 [50] and FSDKaggle2018 [51] for universal sound separation and MUSIC21 [19] for musical instrument separation. In each dataset, evaluation sound mixtures are generated by considering each audio clip as a target source. Subsequently, either 1 or 5 additional audio clips are randomly chosen as interference noise. The mixing process involves combining one target source with one interference noise at a signal-to-noise ratio (SNR) of 0dB, resulting in the creation of 1 or 5 evaluation mixtures for each target source. For label-annotated datasets, we transform the labels into language descriptions by adding the prefix “The sound of ”. For ESC-50, FSDKaggle2018, and MUSIC21, we also randomly choose 10 audio clips for each sound class as query audio samples to evaluate the multi-modal cues queried TSE capability of the proposed approach. These selected query samples are not used to generate evaluation mixtures. In the following section, we will introduce all the evaluation datasets used in our experiments.
@@
TABLE II: Language-queried target sound extraction performance evaluation with SDRs.
@@
TABLE III: Language-queried target sound extraction performance evaluation with CLAPScore.
Modality Polarity CLAPScore $\Delta$CLAPScore CLAPScore $\Delta$CLAPScore CLAPScore $\Delta$CLAPScore CLAPScore $\Delta$CLAPScore CLAPScore $\Delta$CLAPScore
AudioSep [21] Text P 0.365±0.118 0.328±0.194 0.338±0.120 0.274±0.188 0.232±0.110 0.186±0.167 0.317±0.123 0.281±0.184 0.202±0.126 0.167±0.185
CLAPSep Text P 0.369±0.115 0.353±0.175 0.329±0.120 0.254±0.189 0.206±0.138 0.081±0.230 0.312±0.134 0.265±0.219 0.205±0.130 0.192±0.187
N 0.347±0.125 0.341±0.180 0.302±0.132 0.247±0.193 0.155±0.130 0.118±0.160 0.275±0.150 0.269±0.193 0.190±0.135 0.189±0.187
P+N 0.367±0.116 0.365±0.167 0.330±0.120 0.281±0.172 0.218±0.126 0.180±0.165 0.313±0.131 0.313±0.159 0.212±0.124 0.224±0.155
ESC-50 is a labeled collection of environmental audio clips extracted from public field recordings gathered by the Freesound project [52]. The ESC-50 dataset comprises 2,000 5-second-long audio clips belonging to 50 environmental sound classes such as car horn, and pouring water, among others. Each audio clip is assigned to one class label and all the audio clips are sampled at 44.1kHz. For consistency comparison, we downsample the audio clips to 32kHz. Since there is no official test split, we use the whole dataset to create 6,500 evaluation sound mixtures after filtering 500 (10 for each class) audio clips used as query audio samples.
FSDKaggle2018 consists of 11,073 audio files annotated with 41 labels, covering both natural sounds and instrumental music. All the sound samples are gathered from the Freesound project [52] and sampled at 44.1kHz. We resample the audio clips to 32kHz for consistency. To create evaluation sound mixtures, we use the official test split which contains 1,600 audio clips to create 8,000 evaluation sound mixtures. And 410 (10 for each class) query audio samples are randomly sampled from the official train split.
@@ IV-C Evaluation Metrics
Despite SDRs, we also adopt the CLAPScore [54] as another evaluation metric to measure the semantic similarity between the extracted sound source and the text query. The CLAPScore is calculated as:
$\mathrm{CLAPScore}(\hat{{\bm{e}}},{\bm{e}}^{P})=\frac{\hat{{\bm{e}}}^{\top}{\bm{e}}^{P}}{||\hat{{\bm{e}}}||||{\bm{e}}^{P}||},$ (20)
where $\hat{{\bm{e}}}\in\mathbb{R}^{D}$ is obtained by encoding the estimated sound source through the CLAP audio encoder. When the negative queries are accessible, we further propose an improved version of CLAPScore namely $\Delta$CLAPScore, which is defined as:
[…]
This metric further evaluates the ability of target sound extraction systems to suppress non-target sounds, offering a more robust assessment of their performance.
@@ V-A Language-Queried TSE Performance Evaluation
TABLE IV: multi-modal cues queried target sound extraction performance evaluation and SOTA comparison.
@@
The proposed CLAPSep model consistently outperforms other models by a significant margin, except for the MUSIC21 dataset. This anomaly can be traced back to the composition of the AudioCap dataset, on which the CLAPSep trained, contains only 169 music clips among its total of 49,274 clips. Impressively, despite being trained on a considerably smaller dataset than AudioSep, CLAPSep exhibits superior generalizability. This outcome underscores our model’s proficiency in utilizing the extensive sound category knowledge encoded in the pre-trained CLAP model, achieved through integrating the CLAP encoders into both the query and the separation networks to form a tightly coupled system. It is worth noting that CLAPSep supports both positive and negative queries. Its performance with negative prompts is comparable to or even better than that with positive prompts by treating positive and negative queries equally in constructing the condition embedding. Furthermore, the two types of prompts are found to be complementary in the process of extracting target sounds. When CLAPSep is used with both positive and negative prompts, it demonstrates significantly improved performance, particularly on the MUSIC21 dataset. This indicates that models like CLAPSep, which accommodate fine-grained queries, can leverage more complex input queries to effectively extract target sounds, even from sound categories that are rarely represented in the training data. We also visualize the SNRi and SISINRi distribution of CLAPSep and AudioSep across the 5 datasets, as shown in Fig. 4. CLAPSep has higher median scores and is more centrally distributed with lower variance, indicating its better performance and stability.
Alongside the intrusive SDR metric, which reflects the distortion level of the extracted target sound, we also employ CLAPScore [54] to evaluate the semantic similarity between the extracted sound and the query language, as shown in Table III. The results indicate that the proposed method achieves CLAPScore values that are generally comparable to or better than those of the previous state-of-the-art, AudioSep [21], across most datasets. Furthermore, our experiments reveal that the original CLAPScore lacks sensitivity to residual noise in the extracted audio, which limits the apparent performance advantage of our method over AudioSep in this metric. In contrast, the enhanced $\Delta$CLAPScore metric more effectively highlights the superiority of our method, further demonstrating its consistent strengths in both target sound extraction and non-target sound suppression.
@@
V-B Multi-Modal Cues Queried TSE Performance Evaluation
TABLE V: Model performance varying with different numbers of sound sources.
@@ V-B Multi-Modal Cues Queried TSE Performance Evaluation
For comparison, we employ the representative audio-queried TSE model named Universal Source Separation (USS) [9]. Notably, USS is trained on an extensively large-scale dataset, in comparison to ours. The results are presented in Table IV, where “Shots” denotes the number of query audio clips chosen for each sound class. Note that the input query audio clips are converted into fixed-dimensional audio embeddings. When a query has multiple audio samples for a target sound class, the audio embeddings are averaged as one single conditional embedding.
Table IV presents the evaluation results with multi-modal queries. Provided with audio-only queries, the CLAPSep-text model (trained with text-only queries) performs significantly worse than all the others. This underperformance can be attributed to the modality gap phenomenon [42] that substantially hampers the separation decoder trained on a single modality to handle multi-modal queries. In contrast, the model trained with multi-modal queries (denoted as CLAPSep-hybrid) demonstrates remarkable and consistent improvements, affirming the effectiveness of the multi-modal training strategy (as indicated in equation 3) in mitigating the modality gap. Compared with previous SOTA USS, the proposed model achieves better performance on ESC-50 and FSDKaggle2018 when queried by positive queries only. When negative queries are also available, CLAPSep outperforms USS consistently. The query audio embedding can also be augmented with text embedding (referred to as “Audio+Text”). In this case, the proposed CLAPSep model achieves the best performance on most evaluation datasets. This underscores the effectiveness of our proposed method in leveraging multi-modal queries for target sound extraction.
Fig. 5 analyses how the number and duration of query audio samples affect the performance. Firstly, the number of query audio samples, referred to as “Shots,” plays a significant role. When audio samples are selected randomly, a higher number of samples offers richer semantic information, thereby enabling a more accurate representation of the sound class. The duration of query audio samples exhibits a “sweet spot” effect in contrast. Specifically, when the duration of query audio samples reaches 3 seconds, the TSE system demonstrates its optimal performance. This phenomenon suggests that a 3-second clip of query audio can adequately capture the common characteristics of most sound events. Longer duration carves out more detailed but irrelevant characteristics associated with the specific audio clip that can potentially diminish performance.
@@ V-C Zero-Shot Generalizability Evaluation
TABLE VI: Model performance with input SDR varying from -10 to 10 dB.
@@
In this section, we assess the zero-shot capacities of CLAPSep on sound events that are not seen during training. The AudioSet_balanced_train set is employed due to its well-defined classification of event types. When selecting the unseen sound labels, we utilize AudioSet’s official ontology to mark all sound labels under the “Natural sounds” root node as unseen labels, and we filter out all audio samples belonging to these unseen labels from the training set. We use this strategy to avoid semantic correlation between the selected unseen and seen labels.
Subsequently, for each selected unseen category, we evaluate the performance of the TSE system on sound mixtures containing target sounds belonging to that category. The results are shown in Fig. 6 and denoted as “Unseen”. For comparison, we also train CLAPSep without filtering any audio samples, denoted as “Seen”. To mitigate potential label leakage, only positive queries are used for evaluation. The performance gap between “Unseen” and “Seen” is generally minor, indicating that CLAPSep can effectively generalize to sound classes that are unseen during its training phase, which indicates CLAPSep can efficiently inherit the prior knowledge encoded in the multi-modal embedding space of CLAP to identify and extract diverse sound events.
V-D Practical Scenarios Analysis and Evaluation
TABLE VII: Ablation study of language-queried TSE performance.